) A small company borrows $50,000 at 11% compounded monthly. The loan is due in 4 years. How much interest will the company pay? Round to the nearest cent. and please explain how you came up with the answer
Answer by mananth(16946)

Principal = $50,000=p
interest rate = 11% = r
time period = 4 years =n
compounding term = monthly =q
...
A = P(1+[i/q])^nq.
A=50000*(1+(0.11/12))^4*12
...
A=50000*(1+(0.11/12))^48
A = 77,479.90
Interest = A-P
Interest=77,479.90-50,000 = $47,479.90
...
m.ananth@hotmail.ca